A body dropped from a height h with initial velocity 0, strikes the ground with a velocity 3 m/s . Another body of same mass dropped from the same height h with an initial veoocive of 4 m/s. The final velocity of second mass with which it strikes the ground 

Grade:12th pass

1 Answers

reuben
5 years ago
For the 1st object,
u = 0
v = 3m/s 
As the body is dropped from a height,
So a = g = 10m/s2
Using v2 = u2 + 2gh
32 = 02 + 2(10)h
9 = 20h
h = 9/20m
As the second object is also dropped from the same height
So for the second object
h = 9/20m
u = 4m/s
v = ?
Again using v2 = u2 + 2gh
v2 = 42 + 2(10)(9/20)
v2 = 16 + 9
v = root 25
v = 5m/s
The final velocity of second object is 5m/s.